I'm having trouble understanding the question the book is asking. Please read the following passage and questions. The professor does not lecture and the text isn't helping - Thank you
A car and a truck are at a gas station. At 2:00 the truck starts traveling along the highway at a constant speed of 54mph. A quarter of an hour later, at 2:15, the car starts traveling along the same highway with a constant speed of 60mph. At what time will the car catch up with the truck?
Solve this problem algebraically using the following steps.
A) Let Δt be the number of hours the car travels between leaving the gas station and catching up with the truck. Is Δt the answer? If not, how can you get the answer once you found Δt?
I think Δt would be the answer as Δt equals the time between the car and truck...correct?
B) Write an expression for the car's distance from the gas station Δt hours after the car starts. Explain your reasoning.
C) Write an expression for the truck's distance from the gas station Δt hours after the car starts. Explain your reasoning.
D) Write your equation and tell why you set the two sides equal.

Answer by scott8148(6628)

the question is what time (on the clock) does the car catch up

A) Δt is the AMOUNT OF TIME it takes the car to catch the truck
___ it is not the answer
___ answer is 2:15 plus Δt

B) d = 60 * Δt ___ d = r * t

C) d = 54 * (Δt + .25) ___ d = r * t and the truck had a quarter hour head start

D) 60 * Δt = 54 * (Δt + .25) ___ when the car catches the truck, the vehicles had traveled EQUAL distances